OpenStudy (18jonea):

To get to a friend’s house, you walk 6 blocks north, then turn and walk 3 blocks east. What is the displacement of your friend’s house from your starting point? Remember to always use significant digits. 7 blocks 6.7 blocks northeast 6.7 blocks 7 blocks northeast

OpenStudy (18jonea):

@sleepyjess

OpenStudy (18jonea):

@welshfella

OpenStudy (18jonea):

@skullpatrol

OpenStudy (welshfella):

this is a Pythagoras theorem problem 

OpenStudy (welshfella):

X^2 = 6^2 + 3^2

OpenStudy (18jonea):

45

OpenStudy (welshfella):

yes X^2 = 45 so X = ?

OpenStudy (welshfella):

you need the square root of 45 use ur calculator.

OpenStudy (18jonea):

6.7

OpenStudy (18jonea):

@welshfella

OpenStudy (18jonea):

would it be blocks northeast

OpenStudy (18jonea):

7 blocks 6.7 blocks northeast 6.7 blocks 7 blocks northeast these are my options

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

6.7 blocks north east.

OpenStudy (18jonea):

i have one more like this could you check it please?

OpenStudy (welshfella):

ok

OpenStudy (18jonea):

A ball rolls south down the street for 12 meters. It strikes a wall and bounces at a right angle (west) for 9 meters and then comes to a stop. What is the displacement of the ball? Remember to use significant digits. so x^2= 12^2+ 9^2?

OpenStudy (18jonea):

@welshfella

OpenStudy (welshfella):

yes

OpenStudy (18jonea):

225

OpenStudy (18jonea):

so x = 15

OpenStudy (welshfella):

now find sqrt 225

OpenStudy (welshfella):

right and i guess they need the drection

OpenStudy (18jonea):

15 meters southwest?

OpenStudy (welshfella):

yea

OpenStudy (18jonea):

thank you

OpenStudy (welshfella):

yw